# A Reduction-Driven Local Search for the Generalized Independent Set Problem

### v1

@@ A Technical Proofs
Observe that there exists a vertex set $I\in MGIS(G)$ that includes at least one of $u,x,y$. If none of them is included in $I$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(u,v))\geq 0$ In addition, at most one of $u,x,y$ can be included in $I$ because they are permanent neighbors of each other.
Given an arbitrary $I\in MGIS(G)$. When $w(u)\geq\tilde{w}(x)$. If one of $x,y$ is included in $I$,say $x$, then $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)$. In summary, $u$ must be included in some set in $MGIS$.
When $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$. If $y$ is included in $I$, then $u\notin I$ and $nb(I\cup\{u\}\setminus\{y\})\geq nb(I)$. Therefore, there is a vertex set $I^{\prime}\in MGIS(G)$ that $y\notin I^{\prime}$ and exactly one of $u,x$ is included in $I^{\prime}$. We conclude that there exists a vertex set $S\in MGIS(G)$ that $u\in S$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
When $w(u)<\tilde{w}(y)$. $u$ can appear in $I$ iff both $x,y$ are not included in $I$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$ and $y\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$. ∎
@@
When $w(u)\geq\tilde{w}(v)+w^{+}(N(u)\setminus N_{p}[v])$. If $v\in I$, then $u\notin I$. We can replace $N(u)\cap I$ with $u$ to get a larger or equal net benefit because $nb(I\cup\{u\}\setminus N(u))-nb(I)\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\setminus N_{p}[v])\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\cap I)\geq 0$. Therefore, $I\cup\{u\}\setminus N(u)\in MGIS(G)$. Since $v\in N_{p}(u)$, there must be some set in $MGIS(G)$ that does not include $v$.
When $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(u,x))$. If $v\in I$, then $u\notin I$. We can replace $N_{p}(u)\cap I$ with $u$ to get a larger net benefit because $nb(I\cup\{u\}\setminus N_{p}(u))-nb(I)\geq w(u)-\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}(v))+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(u,x))\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\cap I)+\sum_{x\in N_{r}(u)\cap I}\max(0,p(u,x))\geq 0$. Since $v\in N_{p}(u)$, there must be some set in $MGIS(G)$ that does not include $v$. ∎
@@
Assume, for the sake of contradiction, that neither $u$ nor $v$ is included in any set in $MGIS(G)$. If for any $I\in MGIS(G)$, $N(v)\cap I=\varnothing$, then $I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})\geq nb(I)$, a contradiction. Note that $u\notin N(v)\cap I$. If there exists a set $I\in MGIS(G)$ such that $N(v)\cap I\neq\varnothing$, then $w(v)\geq w^{+}(N(v))-max(0,w(u))\geq w^{+}(S)$. $nb(I\cup\{v\}\setminus N(v))\geq nb(I)$, a contradiction.
Therefore, there exists a set in $MGIS(G)$ that includes at least one of $u,v$. We conclude that their common permanent neighbors must not be included in some set in $MGIS(G)$. ∎
@@
Given an arbitrary $I\in MGIS(G)$. If $u$ is included in $I$, then $N_{p}(v)\cap I=\varnothing$. And $I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})-nb(I\setminus\{v\})=w(v)-\sum_{x\in N_{r}(v)}\max(0,p(v,x))\geq 0$. The same argument holds when $v$ is included in $I$. Therefore, we can fold $u$ and $v$ into a vertex $v^{\prime}$. There exists $S\in MGIS(G)$ that $u\in S$ and $v\in S$ iff there exists $S^{\prime}\in MGIS(G^{\prime})$ that $v^{\prime}\in S^{\prime}$. ∎
@@ B Comparative Details on 216 Instances of SET-1
brock200_2_A 200 7416 2460 986 280 986 2 986 $<$1 986 $<$1 986 $<$1
brock400_2_A 400 44847 14939 765 441 765 1 765 $<$1 765 $<$1 765 $<$1
C125.9_A 125 5250 1713 454 $<$1 454 $<$1 454 $<$1 454 $<$1 454 $<$1
C250.9_A 250 20940 7044 581 10 581 $<$1 581 $<$1 581 $<$1 581 $<$1
gen200_p0.9_55_A 200 13434 4176 535 4 535 $<$1 535 $<$1 535 $<$1 535 $<$1
gen400_p0.9_75_A 400 53671 18149 628 82 628 $<$1 628 3 628 $<$1 628 $<$1
hamming8-4_A 256 15580 5284 1094 240 1094 $<$1 1094 $<$1 1094 $<$1 1094 $<$1
keller4_A 171 7057 2378 941 15 941 $<$1 941 $<$1 941 $<$1 941 $<$1
MANN_a27_A 378 53050 17501 533 28 533 $<$1 533 $<$1 533 $<$1 533 $<$1
p_hat300-1_A 300 8239 2694 2308 10800 2680 10800 2744 $<$1 2744 $<$1 2744 $<$1
p_hat300-2_A 300 16403 5525 2076 5182 2076 608 2076 $<$1 2076 $<$1 2076 $<$1
p_hat300-3_A 300 25132 8258 739 100 739 $<$1 739 $<$1 739 $<$1 739 $<$1
brock200_2_B 200 7416 2460 962 276 962 2 962 $<$1 962 $<$1 962 $<$1
brock400_2_B 400 44847 14939 741 445 741 1 741 $<$1 741 $<$1 741 $<$1
C125.9_B 125 5250 1713 437 1 437 $<$1 437 $<$1 437 $<$1 437 $<$1
C250.9_B 250 20940 7044 549 11 549 $<$1 549 $<$1 549 $<$1 549 $<$1
gen200_p0.9_55_B 200 13434 4176 510 5 510 $<$1 510 $<$1 510 $<$1 510 $<$1
gen400_p0.9_75_B 400 53671 18149 595 80 595 $<$1 595 $<$1 595 $<$1 595 $<$1
hamming8-4_B 256 15580 5284 1094 234 1094 $<$1 1094 $<$1 1094 $<$1 1094 $<$1
keller4_B 171 7057 2378 941 15 941 $<$1 941 $<$1 941 $<$1 941 $<$1
MANN_a27_B 378 53050 17501 503 31 503 $<$1 503 $<$1 503 $<$1 503 $<$1
p_hat300-1_B 300 8239 2694 2259 10800 2643 10800 2712 $<$1 2712 $<$1 2712 $<$1
p_hat300-2_B 300 16403 5525 2062 5194 2062 680 2062 $<$1 2062 $<$1 2062 $<$1
p_hat300-3_B 300 25132 8258 713 102 713 $<$1 713 $<$1 713 $<$1 713 $<$1
brock200_2_C 200 7416 2460 932 280 932 2 932 $<$1 932 $<$1 932 $<$1
brock400_2_C 400 44847 14939 698 455 698 2 698 $<$1 698 $<$1 698 $<$1
C125.9_C 125 5250 1713 403 1 403 $<$1 403 $<$1 403 $<$1 403 $<$1
C250.9_C 250 20940 7044 502 12 502 $<$1 502 $<$1 502 $<$1 502 $<$1
gen200_p0.9_55_C 200 13434 4176 467 4 467 $<$1 467 $<$1 467 $<$1 467 $<$1
gen400_p0.9_75_C 400 53671 18149 533 81 533 $<$1 533 $<$1 533 $<$1 533 $<$1
hamming8-4_C 256 15580 5284 1094 229 1094 1 1094 $<$1 1094 $<$1 1094 $<$1
keller4_C 171 7057 2378 941 14 941 $<$1 941 $<$1 941 $<$1 941 $<$1
MANN_a27_C 378 53050 17501 443 30 443 $<$1 443 $<$1 443 $<$1 443 $<$1
p_hat300-1_C 300 8239 2694 2205 10800 2566 10800 2649 $<$1 2649 $<$1 2649 $<$1
p_hat300-2_C 300 16403 5525 2033 5360 2033 778 2033 $<$1 2033 $<$1 2033 $<$1
p_hat300-3_C 300 25132 8258 688 111 688 $<$1 688 $<$1 688 $<$1 688 $<$1
@@
brock200_2_A 200 7416 2460 1489 452 1489 5 1489 $<$1 1489 $<$1 1489 $<$1
brock400_2_A 400 44847 14939 1084 593 1084 2 1084 2 1084 $<$1 1084 $<$1
C125.9_A 125 5250 1713 685 1 685 $<$1 685 $<$1 685 $<$1 685 $<$1
C250.9_A 250 20940 7044 785 12 785 $<$1 785 $<$1 785 $<$1 785 $<$1
gen200_p0.9_55_A 200 13434 4176 778 5 778 $<$1 778 $<$1 778 $<$1 778 $<$1
gen400_p0.9_75_A 400 53671 18149 882 94 882 $<$1 882 $<$1 882 $<$1 882 $<$1
hamming8-4_A 256 15580 5284 1790 331 1790 2 1790 $<$1 1790 $<$1 1790 $<$1
keller4_A 171 7057 2378 1500 23 1500 $<$1 1500 $<$1 1500 $<$1 1500 $<$1
MANN_a27_A 378 53050 17501 683 32 683 $<$1 683 $<$1 683 $<$1 683 $<$1
p_hat300-1_A 300 8239 2694 3981 10800 4182 10800 4674 1 4674 $<$1 4674 $<$1
p_hat300-2_A 300 16403 5525 2994 8743 2994 9402 2994 $<$1 2994 $<$1 2994 $<$1
p_hat300-3_A 300 25132 8258 1180 137 1180 1 1180 $<$1 1180 $<$1 1180 $<$1
brock200_2_B 200 7416 2460 739 456 739 5 739 $<$1 739 $<$1 739 $<$1
brock400_2_B 400 44847 14939 534 611 534 2 534 2 534 $<$1 534 $<$1
C125.9_B 125 5250 1713 335 1 335 $<$1 335 $<$1 335 $<$1 335 $<$1
C250.9_B 250 20940 7044 385 13 385 $<$1 385 $<$1 385 $<$1 385 $<$1
gen200_p0.9_55_B 200 13434 4176 378 5 378 $<$1 378 $<$1 378 $<$1 378 $<$1
gen400_p0.9_75_B 400 53671 18149 432 96 432 $<$1 432 $<$1 432 $<$1 432 $<$1
hamming8-4_B 256 15580 5284 890 333 890 2 890 $<$1 890 $<$1 890 $<$1
keller4_B 171 7057 2378 750 22 750 $<$1 750 $<$1 750 $<$1 750 $<$1
MANN_a27_B 378 53050 17501 333 34 333 $<$1 333 $<$1 333 $<$1 333 $<$1
p_hat300-1_B 300 8239 2694 1981 10800 2082 10800 2324 2 2324 $<$1 2324 $<$1
p_hat300-2_B 300 16403 5525 1494 8947 1494 10800 1494 $<$1 1494 $<$1 1494 $<$1
p_hat300-3_B 300 25132 8258 580 134 580 1 580 1 580 $<$1 580 $<$1
brock200_2_C 200 7416 2460 364 460 364 5 364 $<$1 364 $<$1 364 $<$1
brock400_2_C 400 44847 14939 259 646 259 2 259 1 259 $<$1 259 $<$1
C125.9_C 125 5250 1713 160 1 160 $<$1 160 $<$1 160 $<$1 160 $<$1
C250.9_C 250 20940 7044 185 13 185 $<$1 185 $<$1 185 $<$1 185 $<$1
gen200_p0.9_55_C 200 13434 4176 178 5 178 $<$1 178 $<$1 178 $<$1 178 $<$1
gen400_p0.9_75_C 400 53671 18149 207 98 207 $<$1 207 $<$1 207 $<$1 207 $<$1
hamming8-4_C 256 15580 5284 440 355 440 2 440 $<$1 440 $<$1 440 $<$1
keller4_C 171 7057 2378 375 23 375 $<$1 375 $<$1 375 $<$1 375 $<$1
MANN_a27_C 378 53050 17501 158 33 158 $<$1 158 $<$1 158 $<$1 158 $<$1
p_hat300-1_C 300 8239 2694 981 10800 1032 10800 1149 1 1149 $<$1 1149 $<$1
p_hat300-2_C 300 16403 5525 744 10800 744 10658 744 $<$1 744 $<$1 744 $<$1
p_hat300-3_C 300 25132 8258 280 142 280 $<$1 280 $<$1 280 $<$1 280 $<$1

### v2

@@ A Technical Proofs
Observe that there exists a vertex set $I\in MGIS(G)$ that includes at least one of $u,x,y$. If none of them is included in $I$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(\{u,v\}))\geq 0$ In addition, at most one of $u,x,y$ can be included in $I$ because $\{u,x\}\in E_{p},\{u,y\}\in E_{p},\{x,y\}\in E_{p}$.
Given an arbitrary $I\in MGIS(G)$. When $w(u)\geq\tilde{w}(x)$. If one of $x,y$ is included in $I$,say $x$, then $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)$. In summary, $u$ must be included in at least one set in $MGIS(G)$.
When $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$. If $y$ is included in $I$, then $u\notin I$ and $nb(I\cup\{u\}\setminus\{y\})\geq nb(I)$. Therefore, there is a vertex set $I^{\prime}\in MGIS(G)$ such that $y\notin I^{\prime}$ and exactly one of $u,x$ is included in $I^{\prime}$. We conclude that there exists a vertex set $S\in MGIS(G)$ such that $u\in S$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
When $w(u)<\tilde{w}(y)$. $u$ is in at least one set $I\in MGIS(G)$ iff both $x,y$ are not included in $I$. In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$ and $y\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$. ∎
@@
When $w(u)\geq\tilde{w}(v)+w^{+}(N(u)\setminus N_{p}[v])$. If $v\in I$, then $u\notin I$. We can replace $N(u)\cap I$ with $u$ to get a larger or equal net benefit because $nb(I\cup\{u\}\setminus N(u))-nb(I)\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\setminus N_{p}[v])\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\cap I)\geq 0$. Therefore, $I\cup\{u\}\setminus N(u)\in MGIS(G)$. Since $v\in N_{p}(u)$, there must exist at least one set in $MGIS(G)$ that does not include $v$.
When $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(\{u,x\}))$. If $v\in I$, then $u\notin I$. We can replace $N_{p}(u)\cap I$ with $u$ to get a larger net benefit because $nb(I\cup\{u\}\setminus N_{p}(u))-nb(I)\geq w(u)-\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}(v))+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(\{u,x\}))\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\cap I)+\sum_{x\in N_{r}(u)\cap I}\max(0,p(\{u,x\}))\geq 0$. Since $v\in N_{p}(u)$, there must exist a set in $MGIS(G)$ that does not include $v$. ∎
@@
Assume, for the sake of contradiction, that neither $u$ nor $v$ is included in any set in $MGIS(G)$. If for any $I\in MGIS(G)$, $N(v)\cap I=\varnothing$, then $I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})\geq nb(I)$ without containing permanent edges, a contradiction. Note that $u\notin N(v)\cap I$. If there exists a set $I\in MGIS(G)$ such that $N(v)\cap I\neq\varnothing$, then $w(v)\geq w^{+}(N(v))-max(0,w(u))\geq w^{+}(S)$. $nb(I\cup\{v\}\setminus N(v))\geq nb(I)$ without containing permanent edges, again a contradiction. Therefore, there exists a set in $MGIS(G)$ that includes at least one of $u,v$. We conclude that $N_{p}(u)\cap N_{p}(v)$ must not be included in at least one set in $MGIS(G)$. ∎
@@
Given an arbitrary $I\in MGIS(G)$. If $u$ is included in $I$, then $N_{p}(v)\cap I=\varnothing$. $I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})-nb(I\setminus\{v\})=w(v)-\sum_{x\in N_{r}(v)}\max(0,p(\{v,x\}))\geq 0$. The same argument holds when $v$ is included in $I$. Therefore, we can fold $u$ and $v$ into a vertex $v^{\prime}$. There exists $S\in MGIS(G)$ such that $u\in S$ and $v\in S$ iff there exists $S^{\prime}\in MGIS(G^{\prime})$ such that $v^{\prime}\in S^{\prime}$. ∎
@@ B Comparative Details on 216 Instances of SET-1
brock200_2_A 200 7,416 2,460 986 280 986 2 986 <<1 986 <<1 986 <<1
brock400_2_A 400 44,847 14,939 765 441 765 1 765 <<1 765 <<1 765 <<1
C125.9_A 125 5,250 1,713 454 <<1 454 <<1 454 <<1 454 <<1 454 <<1
C250.9_A 250 20,940 7,044 581 10 581 <<1 581 <<1 581 <<1 581 <<1
gen200_p0.9_55_A 200 13,434 4,176 535 4 535 <<1 535 <<1 535 <<1 535 <<1
gen400_p0.9_75_A 400 53,671 18,149 628 82 628 <<1 628 3 628 <<1 628 <<1
hamming8-4_A 256 15,580 5,284 1,094 240 1,094 <<1 1,094 <<1 1,094 <<1 1,094 <<1
keller4_A 171 7,057 2,378 941 15 941 <<1 941 <<1 941 <<1 941 <<1
MANN_a27_A 378 53,050 17,501 533 28 533 <<1 533 <<1 533 <<1 533 <<1
p_hat300-1_A 300 8,239 2,694 2,308 10,800 2,680 10,800 2,744 <<1 2,744 <<1 2,744 <<1
p_hat300-2_A 300 16,403 5,525 2,076 5,182 2,076 608 2,076 <<1 2,076 <<1 2,076 <<1
p_hat300-3_A 300 25,132 8,258 739 100 739 <<1 739 <<1 739 <<1 739 <<1
brock200_2_B 200 7,416 2,460 962 276 962 2 962 <<1 962 <<1 962 <<1
brock400_2_B 400 44,847 14,939 741 445 741 1 741 <<1 741 <<1 741 <<1
C125.9_B 125 5,250 1,713 437 1 437 <<1 437 <<1 437 <<1 437 <<1
C250.9_B 250 20,940 7,044 549 11 549 <<1 549 <<1 549 <<1 549 <<1
gen200_p0.9_55_B 200 13,434 4,176 510 5 510 <<1 510 <<1 510 <<1 510 <<1
gen400_p0.9_75_B 400 53,671 18,149 595 80 595 <<1 595 <<1 595 <<1 595 <<1
hamming8-4_B 256 15,580 5,284 1,094 234 1,094 <<1 1,094 <<1 1,094 <<1 1,094 <<1
keller4_B 171 7,057 2,378 941 15 941 <<1 941 <<1 941 <<1 941 <<1
MANN_a27_B 378 53,050 17,501 503 31 503 <<1 503 <<1 503 <<1 503 <<1
p_hat300-1_B 300 8,239 2,694 2,259 10,800 2,643 10,800 2,712 <<1 2,712 <<1 2,712 <<1
p_hat300-2_B 300 16,403 5,525 2,062 5,194 2,062 680 2,062 <<1 2,062 <<1 2,062 <<1
p_hat300-3_B 300 25,132 8,258 713 102 713 <<1 713 <<1 713 <<1 713 <<1
brock200_2_C 200 7,416 2,460 932 280 932 2 932 <<1 932 <<1 932 <<1
brock400_2_C 400 44,847 14,939 698 455 698 2 698 <<1 698 <<1 698 <<1
C125.9_C 125 5,250 1,713 403 1 403 <<1 403 <<1 403 <<1 403 <<1
C250.9_C 250 20,940 7,044 502 12 502 <<1 502 <<1 502 <<1 502 <<1
gen200_p0.9_55_C 200 13,434 4,176 467 4 467 <<1 467 <<1 467 <<1 467 <<1
gen400_p0.9_75_C 400 53,671 18,149 533 81 533 <<1 533 <<1 533 <<1 533 <<1
hamming8-4_C 256 15,580 5,284 1,094 229 1,094 1 1,094 <<1 1,094 <<1 1,094 <<1
keller4_C 171 7,057 2,378 941 14 941 <<1 941 <<1 941 <<1 941 <<1
MANN_a27_C 378 53,050 17,501 443 30 443 <<1 443 <<1 443 <<1 443 <<1
p_hat300-1_C 300 8,239 2,694 2,205 10,800 2,566 10,800 2,649 <<1 2,649 <<1 2,649 <<1
p_hat300-2_C 300 16,403 5,525 2,033 5,360 2,033 778 2,033 <<1 2,033 <<1 2,033 <<1
p_hat300-3_C 300 25,132 8,258 688 111 688 <<1 688 <<1 688 <<1 688 <<1
@@
brock200_2_A 200 7,416 2,460 1,489 452 1,489 5 1,489 <<1 1,489 <<1 1,489 <<1
brock400_2_A 400 44,847 14,939 1,084 593 1,084 2 1,084 2 1,084 <<1 1,084 <<1
C125.9_A 125 5,250 1,713 685 1 685 <<1 685 <<1 685 <<1 685 <<1
C250.9_A 250 20,940 7,044 785 12 785 <<1 785 <<1 785 <<1 785 <<1
gen200_p0.9_55_A 200 13,434 4,176 778 5 778 <<1 778 <<1 778 <<1 778 <<1
gen400_p0.9_75_A 400 53,671 18,149 882 94 882 <<1 882 <<1 882 <<1 882 <<1
hamming8-4_A 256 15,580 5,284 1,790 331 1,790 2 1,790 <<1 1,790 <<1 1,790 <<1
keller4_A 171 7,057 2,378 1,500 23 1,500 <<1 1,500 <<1 1,500 <<1 1,500 <<1
MANN_a27_A 378 53,050 17,501 683 32 683 <<1 683 <<1 683 <<1 683 <<1
p_hat300-1_A 300 8,239 2,694 3,981 10,800 4,182 10,800 4,674 1 4,674 <<1 4,674 <<1
p_hat300-2_A 300 16,403 5,525 2,994 8,743 2,994 9,402 2,994 <<1 2,994 <<1 2,994 <<1
p_hat300-3_A 300 25,132 8,258 1,180 137 1,180 1 1,180 <<1 1,180 <<1 1,180 <<1
brock200_2_B 200 7,416 2,460 739 456 739 5 739 <<1 739 <<1 739 <<1
brock400_2_B 400 44,847 14,939 534 611 534 2 534 2 534 <<1 534 <<1
C125.9_B 125 5,250 1,713 335 1 335 <<1 335 <<1 335 <<1 335 <<1
C250.9_B 250 20,940 7,044 385 13 385 <<1 385 <<1 385 <<1 385 <<1
gen200_p0.9_55_B 200 13,434 4,176 378 5 378 <<1 378 <<1 378 <<1 378 <<1
gen400_p0.9_75_B 400 53,671 18,149 432 96 432 <<1 432 <<1 432 <<1 432 <<1
hamming8-4_B 256 15,580 5,284 890 333 890 2 890 <<1 890 <<1 890 <<1
keller4_B 171 7,057 2,378 750 22 750 <<1 750 <<1 750 <<1 750 <<1
MANN_a27_B 378 53,050 17,501 333 34 333 <<1 333 <<1 333 <<1 333 <<1
p_hat300-1_B 300 8,239 2,694 1,981 10,800 2,082 10,800 2,324 2 2,324 <<1 2,324 <<1
p_hat300-2_B 300 16,403 5,525 1,494 8,947 1,494 10,800 1,494 <<1 1,494 <<1 1,494 <<1
p_hat300-3_B 300 25,132 8,258 580 134 580 1 580 1 580 <<1 580 <<1
brock200_2_C 200 7,416 2,460 364 460 364 5 364 <<1 364 <<1 364 <<1
brock400_2_C 400 44,847 14,939 259 646 259 2 259 1 259 <<1 259 <<1
C125.9_C 125 5,250 1,713 160 1 160 <<1 160 <<1 160 <<1 160 <<1
C250.9_C 250 20,940 7,044 185 13 185 <<1 185 <<1 185 <<1 185 <<1
gen200_p0.9_55_C 200 13,434 4,176 178 5 178 <<1 178 <<1 178 <<1 178 <<1
gen400_p0.9_75_C 400 53,671 18,149 207 98 207 <<1 207 <<1 207 <<1 207 <<1
hamming8-4_C 256 15,580 5,284 440 355 440 2 440 <<1 440 <<1 440 <<1
keller4_C 171 7,057 2,378 375 23 375 <<1 375 <<1 375 <<1 375 <<1
MANN_a27_C 378 53,050 17,501 158 33 158 <<1 158 <<1 158 <<1 158 <<1
p_hat300-1_C 300 8,239 2,694 981 10,800 1,032 10,800 1,149 1 1,149 <<1 1,149 <<1
p_hat300-2_C 300 16,403 5,525 744 10,800 744 10,658 744 <<1 744 <<1 744 <<1
p_hat300-3_C 300 25,132 8,258 280 142 280 <<1 280 <<1 280 <<1 280 <<1